OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle? The square root of 149 51 The square root of 51 3

OpenStudy (hoblos):

 use the theorem: h^2 = a^2 + b^2 so 10^2 = 7^2 + x^2 x^2 = 100-49=51 thus x=square root of 51

OpenStudy (anonymous):

okay thanks

OpenStudy (hoblos):

my pleasure :D